OpenStudy (anonymous):

Please can anybody help me solve? Lamar invested his savings in two investment funds. The amount he invested in fund A was $2000.00 less than the amount he invested in fund B . Fund A returned a 8% profit and fund B returned a 4% profit. How much did he invest in fund A if the total profit from the two funds together was $800.00. Amount invested in fund A =$

OpenStudy (anonymous):

(x-2000)+((x-2000)*0.08)+x+0.04x=800

OpenStudy (anonymous):

(x-2000)+((x-2000)*0.08)+1.04x=800

OpenStudy (anonymous):

(2.04x-2000)+((x-2000)*0.08)=800

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

(2.04x-2000)+(0.08x-160)=800

OpenStudy (anonymous):

2.12x-2160=800 Glad to help.

OpenStudy (anonymous):

You going to close this?